find angle A if it supplements angle B which is equal to 70-x
Answer by RAY100(1637)

A + B = 180,,,,def supplementary
.
B= (70-x),,,,given
.
A +(70-x) =180
.
A + (70-x) = 180
.
A = 180 -70 +x
.
A= 110 +x
.
check, x=10, (70-x) =60,,,A=120,,,answer,,,110+10=120,,,,ok